OpenStudy (anonymous):

Evaluate. 9!/7! a) 63 b) 72 c) 81

OpenStudy (anonymous):

Cancel.

jimthompson5910 (jim_thompson5910):

You can take advantage of the fact that 9! = 9*8*7!

jimthompson5910 (jim_thompson5910):

Or you can expand everything out to get 9! = 9*8*7*6*5*4*3*2*1 and 7! = 7*6*5*4*3*2*1

OpenStudy (anonymous):

Canceling is more efficient. Use the observation made via the definition of the factorial by Jim.

OpenStudy (anonymous):

which is ??

jimthompson5910 (jim_thompson5910):

9! ---- 7! becomes 9*8*7! ------- 7! and cancels to 9*8 ----- 1

jimthompson5910 (jim_thompson5910):

or just 9*8

OpenStudy (anonymous):

@burakb, please do not post full answers. It's against the Code of Conduct.

OpenStudy (anonymous):

Alright 9!=9*8*7*6*5*4*3*2*1 and 7!=7*6*5*4*3*2*1 so \[9!\div7!=(9*8*7*6*5*4*3*2*1*)\div(7*6*5*4*3*2*1)=9*8=72\]